# The decomposition of primes in nonabelian extensions of Heisenberg type and an analogue of Euler’s criterion

### v2

@@ 4. $\ell\geq 3$
Now suppose that the conjugacy class of Frobenius element $\operatorname{Frob}_{(t-a),\mathcal{R}^{(\ell)}}$ is the identity element. Say the restriction of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(\ell)}}$ to $L=k(t)(t^{1/\ell})$ as $\sigma$. Then, $\sigma$ is the identity element of $\operatorname{Gal}(L/k(t))$ from our assumption. This contradicts with $\genfrac{(}{)}{}{}{a}{p}_{\ell}=1$. Indeed, its induced map on $\mathcal{O}_{L}/(t-a)\to\mathcal{O}_{L}/(t-a)$ should be the identity map. On the other hand, $\mathcal{O}_{L}\cong k[t^{1/\ell}]$ implies $\mathcal{O}_{L}/(t-a)\cong k[t]/(t^{\ell}-a)$. From the fact that $L/k(t)$ is an abelian extenion, $\operatorname{Frob}_{(t-a),L}(t^{1/\ell})=t^{p/\ell}=a^{p-1/\ell}t^{1/\ell}=\genfrac{(}{)}{}{}{a}{p}_{\ell}t^{1/\ell}$. Thus, if $\genfrac{(}{)}{}{}{a}{p}_{\ell}\neq 1$, $\operatorname{Frob}_{(t-a),\mathcal{R}^{(\ell)}}$ cannot be the identity map. This completes the argument for the first case.
@@
First, we show that the ring of integers of $\mathcal{K}^{(\ell)}$ is given by $k[t][t^{1/\ell},(1-t)^{1/\ell}]$. Next, we calculate the relative discriminant $\Delta_{\mathcal{R}^{(\ell)}/\mathcal{K}^{(\ell)}}$. Finally, we find the integral basis of $\mathcal{R}^{(\ell)}/\mathcal{K}^{(\ell)}$. The basis we have found are almost Galois invariant in the sense that the line spanned by them are invariant under the Galois action.
@@
Let $\zeta=\zeta_{\ell}$, $I_{n}=(t^{1/\ell}-\zeta^{n},(1-t)^{1/\ell})$. First, we observe that $I_{n}$ is a prime ideal in $\mathcal{O}_{\mathcal{K}^{(\ell)}}$. Indeed, $\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]/I_{n}\cong\mathbb{F}_{p}$.
Next, we first show that $(t^{1/\ell}-\zeta^{n})\supseteq I_{n}^{\ell}$ holds. We need to show $\prod_{i=1}^{\ell}(a_{i}(t^{1/\ell}-\zeta^{n})+b_{i}(1-t)^{1/\ell})\in(t^{1/\ell}-\zeta^{n})$ for all $a_{i},b_{i}\in\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]$. After expanding the product, all the terms except $\prod_{i=1}^{\ell}b_{i}(1-t)^{1/\ell}=(\prod_{i=1}^{\ell}b_{i})(1-t)$ is already divisible by $t^{1/\ell}-\zeta^{n}$. Also, since $(1-t)=\prod_{i=1}^{\ell}(1-\zeta^{i}t^{1/\ell})$, the product $\prod_{i=1}^{\ell}(a_{i}(t^{1/\ell}-\zeta^{n})+b_{i}(1-t)^{1/\ell})$ is an element of the ideal $(t^{1/\ell}-\zeta^{n})$.
Finally, we show that $(t^{1/\ell}-\zeta^{n})\subseteq I_{n}^{\ell}$ holds. We only need to verify $t^{1/\ell}-\zeta^{n}\in I_{n}^{\ell}$. We claim that $t^{1/\ell}-\zeta^{n}=a(t^{1/\ell}-\zeta^{n})^{\ell}+b(1-t)$ for $a,b\in\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]$. Substituting $x=t^{1/\ell}$, we observe that $(x-\zeta^{n})^{n}$ and $(1-x^{\ell})$ has greatest common divisor $x-\zeta^{n}$ in $\mathbb{F}_{p}[x]$. Hence, by Bezout’s identity, we get $a,b\in\mathbb{F}_{p}[x]\subset\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]$ with $t^{1/\ell}-\zeta^{n}=a(t^{1/\ell}-\zeta^{n})^{\ell}+b(1-t)$. ∎
@@
Now, the discriminant includes the ideal generated by the square of determinants of matrices of the form $(\zeta_{\ell}^{i}\epsilon_{\ell,n}(t)^{j/\ell})_{ij}$ for $0\leq n<\ell$. The determinants are given by $\left(\prod_{i<j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)\epsilon_{\ell,n}(t)^{(\ell-1)/2}$. Also, note that $(\epsilon_{\ell,n}(t))=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell})^{i}=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell},(1-t)^{1/\ell})^{i\ell}$ as an ideal. By the lemma above, greatest common divisor of these ideals for $0\leq n\leq\ell-1$ is the unit ideal, which shows that the relative discriminant is the unit ideal. ∎
@@
Let $A_{1},\dots,A_{\ell}$ be $n\times n$ matrices. Put
@@
Then, $\det(A)=\prod_{i}\det(A_{i})\times\left(\prod_{i\neq j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)^{n}$.
@@
From the formula for determinant of the Vandermonde matrix, $\det B=\left(\prod_{i\neq j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)$. Now, consider the series of row operations $\Set{R_{i}}$ that transforms $B$ into diagonal matrix $B^{\prime}$. Then, the product of diagonal elements is given by $\det B$. Now, apply the same row action on $A$ considering it as a $\ell\times\ell$ matrix with $n\times n$ block matrix as its entry. Then, we get the block diagonal matrix with diagonal entries $B^{\prime}_{ii}A_{i}$. Hence, $\det A=\prod_{i}\det(B_{ii}^{\prime}A_{i})=(\det B)^{n}\times\prod_{i}\det(A_{i})$
@@
First, $\det A_{0}$ is exactly given when we calculated the discriminant of $\mathcal{K}^{(\ell)}/\mathbb{F}_{p}(t)$ which is $C\times(t(1-t))^{\frac{\ell(\ell-1)}{2}}$. Next, we calculate $\det A_{k}$ for $k>0$.
Now, fixing $y,j$, we can express $A_{k}$ by
@@
To prove the lemma, we show that $\mathbb{F}_{p}[t][\alpha_{i,j}^{k}]/(t^{1/\ell}-\zeta_{\ell}^{i}x,(1-t)^{1/\ell}-\zeta_{\ell}^{j}y)$ is isomorphic to $\mathbb{F}_{p}[\epsilon_{\ell}(t)^{1/\ell}]/(\epsilon_{\ell}(t)-\epsilon_{\ell}(a))$.

### v3

@@ 4. $\ell\geq 3$
Now suppose that the conjugacy class of Frobenius element $\operatorname{Frob}_{(t-a),\mathcal{R}^{(\ell)}}$ is the identity element. Denote the restriction of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(\ell)}}$ to $L=k(t)(t^{1/\ell})$ by $\sigma$. Then, $\sigma$ is the identity element of $\operatorname{Gal}(L/k(t))$ from our assumption. This contradicts with $\genfrac{(}{)}{}{}{a}{p}_{\ell}=1$. Indeed, its induced map on $\mathcal{O}_{L}/(t-a)\to\mathcal{O}_{L}/(t-a)$ should be the identity map. On the other hand, $\mathcal{O}_{L}\cong k[t^{1/\ell}]$ implies $\mathcal{O}_{L}/(t-a)\cong k[t]/(t^{\ell}-a)$. From the fact that $L/k(t)$ is an abelian extenion, $\operatorname{Frob}_{(t-a),L}(t^{1/\ell})=t^{p/\ell}=a^{p-1/\ell}t^{1/\ell}=\genfrac{(}{)}{}{}{a}{p}_{\ell}t^{1/\ell}$. Thus, if $\genfrac{(}{)}{}{}{a}{p}_{\ell}\neq 1$, $\operatorname{Frob}_{(t-a),\mathcal{R}^{(\ell)}}$ cannot be the identity map. This completes the argument for the first case.
@@
First, we show that the ring of integers of $\mathcal{K}^{(\ell)}$ is given by $k[t][t^{1/\ell},(1-t)^{1/\ell}]$. Next, we calculate the relative discriminant $\Delta_{\mathcal{R}^{(\ell)}/k(t)}$. Finally, we find the integral basis of $\mathcal{R}^{(\ell)}/k(t)$. The basis we have found is almost Galois invariant in the sense that the line spanned by them are invariant under the Galois action.
@@
To simplify the notation, put $\zeta=\zeta_{\ell}$ and $I_{n}=(t^{1/\ell}-\zeta^{n},(1-t)^{1/\ell})$. First, we observe that $I_{n}$ is a prime ideal in $\mathcal{O}_{\mathcal{K}^{(\ell)}}$. Indeed, $\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]/I_{n}\cong\mathbb{F}_{p}$.
Next, we first show that $(t^{1/\ell}-\zeta^{n})\supseteq I_{n}^{\ell}$ holds. We need to show $\prod_{i=1}^{\ell}(a_{i}(t^{1/\ell}-\zeta^{n})+b_{i}(1-t)^{1/\ell})\in(t^{1/\ell}-\zeta^{n})$ for all $a_{i},b_{i}\in\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]$. After expanding the product, all the terms except $\prod_{i=1}^{\ell}b_{i}(1-t)^{1/\ell}=(\prod_{i=1}^{\ell}b_{i})(1-t)$ is already divisible by $t^{1/\ell}-\zeta^{n}$. Also, since $(1-t)=\prod_{i=1}^{\ell}(\zeta^{i}-t^{1/\ell})$, the product $\prod_{i=1}^{\ell}(a_{i}(t^{1/\ell}-\zeta^{n})+b_{i}(1-t)^{1/\ell})$ is an element of the ideal $(t^{1/\ell}-\zeta^{n})$.
Finally, we show that $(t^{1/\ell}-\zeta^{n})\subseteq I_{n}^{\ell}$ holds. We only need to verify $t^{1/\ell}-\zeta^{n}\in I_{n}^{\ell}$. We claim that there exists $a,b\in\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]$ satisfying $t^{1/\ell}-\zeta^{n}=a(t^{1/\ell}-\zeta^{n})^{\ell}+b(1-t)$. Substituting $x=t^{1/\ell}$, we observe that $(x-\zeta^{n})^{n}$ and $(1-x^{\ell})$ has greatest common divisor $x-\zeta^{n}$ in $\mathbb{F}_{p}[x]$. Hence, by Bezout’s identity, we get $a,b\in\mathbb{F}_{p}[x]\subset\mathbb{F}_{p}[t][t^{1/\ell},(1-t)^{1/\ell}]$ with $t^{1/\ell}-\zeta^{n}=a(t^{1/\ell}-\zeta^{n})^{\ell}+b(1-t)$. ∎
@@
Now, the discriminant includes the ideal generated by the square of determinants of matrices of the form $(\zeta_{\ell}^{i}\epsilon_{\ell,n}(t)^{j/\ell})_{ij}$ for $0\leq n<\ell$. The determinants are given by $\left(\prod_{i<j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)\epsilon_{\ell,n}(t)^{(\ell-1)/2}$. Also, note that $(\epsilon_{\ell,n}(t))=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell})^{i}=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell},(1-t)^{1/\ell})^{i\ell}$ as an ideal by the lemma above. Thus, greatest common divisor of these ideals for $0\leq n\leq\ell-1$ is the unit ideal, which shows that the relative discriminant is the unit ideal. ∎
@@
Let $A_{1},\dots,A_{\ell}$ be $n\times n$ matrices. Put $\ell n\times\ell n$ matrix
@@
Then, $\det(A)=\prod_{i=1}^{\ell}\det(A_{i})\times\left(\prod_{i\neq j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)^{n}$.
@@
From the formula for determinant of the Vandermonde matrix, $\det B=\left(\prod_{i\neq j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)$. Now, consider the series of row operations $\Set{R_{i}}$ that transforms $B$ into diagonal matrix $B^{\prime}$. Then, the product of diagonal elements is given by $\det B$. Now, apply the same row action on $A$ considering it as a $\ell\times\ell$ matrix with $n\times n$ block matrix as its entry. Then, we get the block diagonal matrix with diagonal entries $B^{\prime}_{ii}A_{i}$. Hence, $\det A=\prod_{i=1}^{\ell}\det(B_{ii}^{\prime}A_{i})=(\det B)^{n}\times\prod_{i=1}^{\ell}\det(A_{i})$
@@
First, $\det A_{0}$ is exactly the discriminant of $\mathcal{K}^{(\ell)}/\mathbb{F}_{p}(t)$ which is $C\times(t(1-t))^{\frac{\ell(\ell-1)}{2}}$. Next, we calculate $\det A_{k}$ for $k>0$.
Now, varying $y,j$, we can express $A_{k}$ by
@@
To prove the lemma, we show that
$\mathbb{F}_{p}[t][\alpha_{i,j}^{k}]/(t^{1/\ell}-\zeta_{\ell}^{i}x,(1-t)^{1/\ell}-\zeta_{\ell}^{j}y)$
is isomorphic to $\mathbb{F}_{p}[\epsilon_{\ell}(t)^{1/\ell}]/(\epsilon_{\ell}(t)-\epsilon_{\ell}(a))$.